A student must answer four of the six essay questions on a history test. how many different selections are possible?
Answer by jim_thompson5910(35256)

There are 6*5*4*3 = 360 different ways to choose the questions where order matters.

Since order does not matter, we divide by 4! = 4*3*2*1 = 24 to get 360/24 = 15

There are 15 different ways to do this.